how many different ways can the position of first base, second base, third base and short stop be filled with a baseball team that consists of 14 players?
Answer by Edwin McCravy(20054)

14 Position 4 = permutations of 14 things taken 4 at a time = P(14,4) = 14×13×12×11 = 24,024